OpenStudy (anonymous):

What is the radius of a circle with a circumference of 14.13 meters? Use 3.14 for pi and show your work.

OpenStudy (amistre64):

we simply need to find a formula that relates circumference to radius right?

OpenStudy (anonymous):

Yes.

OpenStudy (amistre64):

I beleive we can do it with C = 2pi r

OpenStudy (anonymous):

c=2pi.r 14.13=2.(3.14).r r= 2.25

OpenStudy (amistre64):

if we sove this for "r" we get: C/2pi = r , and fill in the information we have :)

OpenStudy (amistre64):

\[\frac{14.13}{2*3.14}=r\] and math it thru

OpenStudy (anonymous):

So the answer is....

OpenStudy (amistre64):

well, that part I want you to do, and I can check your work if you want :) raheen might be accurate enough in it tho

OpenStudy (anonymous):

OK.

OpenStudy (amistre64):

the problem itself is to get you to be able to use the formula and manipulate it to find the answer; after that, its just math .. to calculate by hand or by machine ... http://www.wolframalpha.com/input/?i=14.13%2F%282*3.14%29